Eliminate the parameter from the parametric equations x=1/2t+6 and
y = cos t - sin t.

Answer by fractalier(6550)

Let's solve this
x=1/2t+6
for t and then plug it in to the other...
x - 6 = (1/2)t
t = 2(x-6) = 2x - 12
Then
y = cos(2x - 12) - sin(2x - 12)